A random sample of size 4 is taken with replacement from a population with population mean of 12 and population variance of 8. What is the variance of the sampling distribution.


1
Expert's answer
2021-05-07T10:32:14-0400

Given, Population mean "\mu=12" , population variance "\sigma^2=8"

sample size "n=4"


Sample standrad deviation "s= \sqrt{\sigma^2}\times \sqrt{n}=\sqrt{8}\times \sqrt{4}=\sqrt{32}"


Variance of sampling Distribution "s^2=(\sqrt{32})^2=32"